OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x3 - 6x2 + 9x

OpenStudy (***[isuru]***):

factorize it..this time it's ur turn.. :) first take out a common factor.. u can take out "x" as the common factor ... wt will u get then ?

OpenStudy (anonymous):

3-6(2)+9????? omg I dunno

OpenStudy (***[isuru]***):

x [ x^2 - 6x + 9 ]

OpenStudy (***[isuru]***):

now u can factorize the other parts x [ x^2 - 6x + 9 ] = x [ x^2 -3x - 3x + 9 ] = x[ x ( x -3 ) - 3( x -3 ) ] = x [ ( x-3) ( x-3) ]

OpenStudy (anonymous):

ooooohhh I c I didn't put in the exponents the right way, hahaha its supposed to be x^3-6x^2+9x

OpenStudy (***[isuru]***):

so... if the function is equal to zero x = 0 or x= 3 which means zeros r 0 and 3

OpenStudy (anonymous):

yup.....ur so right! :D